What are the measures of two complementary angles if the difference in the measures of the two angles is 12?
Answer by nerdybill(5404)

What are the measures of two complementary angles if the difference in the measures of the two angles is 12?
.
if two angles are complementary, then the sum is 90 deg.
.
Let x = larger of the two angles
then
90-x = smaller of the two
.
x - (90-x) = 12
x - 90 + x = 12
2x - 90 = 12
2x = 102
x = 51 deg (larger angle)
.
smaller angle:
90-x = 90-51 = 39 deg